Circular No. 7471

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 S4 (LINEAR) Z. Sekanina, Jet Propulsion Laboratory, reports: “The unusually large

nongravitational forces found by B. G. Marsden (MPEC 2000-O07) suggest that comet C/1999 S4 was a trailing fragment of a more massive comet that has been moving in the same orbit, arrived at perihelion long (centuries?) ago but—not surprisingly—was missed. Trailing fragments of known comet pairs have a tendency to sudden disintegration (e.g., Sekanina 1997, A.Ap. 318, L5). If much of the comet’s mass did indeed dissipate into a cloud of dust in the recent event, as suggested by M. R. Kidger (IAUC 7467) and others, the total mass involved could be estimated by further monitoring the tail. Experience with the past initially bright comets that later became headless and disappeared shows that a narrow, bandlike tail—a developing synchronic formation—should survive the head by several weeks or even longer (Sekanina 1984, Icarus 58, 81). A very preliminary analysis suggests that the event may have begun as early as July 23.6 UT and involved submillimeter-sized and larger dust (repulsive accelerations β ≤ 0.024 of the solar attraction). The position angle and approximate length of this tail feature are then predicted to reach: July 30.0 UT, 90^{o}, 2′; Aug. 4.0, 98^{o}, 4′; 9.0, 102^{o}, 7′; 14.0, 104^{o}, 10′; 19.0, 105^{o}, 12′; 24.0, 106^{o}, 15′; 29.0, 106^{o}, 17′. Especially toward the end of this period, the predicted length probably is a crude upper bound. If no such tail persists, the comet’s upper mass limit should be tightly constrained, or the amount of dust lost in the event did not represent a substantial fraction of the total mass.”

Total visual magnitudes: July 25.90 UT, 6.6 (M. Lehký, Hradec Králové, Czech Republic, 10 × 80 binoculars); 26.87, 6.5 (J. Kysely, Vlasim, Czech Republic, 7 × 50 binoculars); 27.86, 6.6 (M. Reszelski, Szamotuly, Poland, 20 × 60 binoculars); 28.86, 7.0 (M. V. Zanotta, Alpe di Ponna, Italy, 7 × 42 binoculars); 29.48, 7.4 (Y. Nagai, Yamanashi, Japan, 20 × 100 binoculars); 30.47, 7.9 (S. Yoshida, Ibaraki, Japan, 0.25-m reflector).

SUPERNOVA 2000cx IN NGC 524 D. C. Leonard, A. V. Filippenko, R. Chornock and W. D. Li, Univer-

sity of California at Berkeley, report: “Spectropolarimetry (range 430–690 nm) obtained on July 28 with the Shane 3-m reflector, Lick Observatory, indicates continuum polarization of ∼ 0.5 percent. A strong polarization modulation (∼ 0.3 percent) across the Si II (635.5 nm) P-Cyg feature implies significant polarization intrinsic to the supernova itself, possibly resulting from an aspherical scattering atmosphere (e.g., Hoeflich 1991, A.Ap. 246, 481). According to photometry by KAIT (see IAUC 6627, 7126), SN 2000cx reached maximum B brightness (preliminary B = 13.3) on July 27.

2000 July 30  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Brian G. Marsden